Fogerty Company makes two products—titanium Hubs and Sprockets. Data regarding the two products follow:

Direct
Labor-Hours per Unit
Annual
Production
Hubs 0.80 16,000 units
Sprockets 0.40 47,000 units

Additional information about the company follows:

  1. Hubs require $27 in direct materials per unit, and Sprockets require $13.

  2. The direct labor wage rate is $10 per hour.

  3. Hubs require special equipment and are more complex to manufacture than Sprockets.

  4. The ABC system has the following activity cost pools:

Estimated Activity
Activity Cost Pool (Activity Measure) Overhead Cost Hubs Sprockets Total
Machine setups (number of setups) $ 28,080 130 104 234
Special processing (machine-hours) $ 144,000 4,800 0 4,800
General factory (organization-sustaining) $ 125,600 NA NA NA

Required:

1. Compute the activity rate for each activity cost pool.

Activity Cost Pool Activity Rate
Machine setups per setup
Special processing per MH

2. Determine the unit product cost of each product according to the ABC system.

Hubs Sprockets
Direct materials
Direct labor
Overhead
Unit cost
Answer #1

1) Activity rate

Activity cost pool Activity rate
Machine setups 28080/234 = 120 per setup
Special processing 144000/4800 = 30 per MH

2) Calculate unit product cost

Hubs Sprockets
Direct material 27 13
Direct labor 10*.80 = 8 10*.40 = 4
Overhead (130*120+144000/16000) = 9.975 (104*120/47000) = 0.266
Unit cost 44.975 17.266